OpenStudy (anonymous):

1. If I know the lengths of two sides of a right triangle, how do I find the third? 2. Could I find the two missing side lengths of a right triangle if I only know one side length and one angle measure (other than the 90 degree angle)? 3. Could I find the two missing angle measures if I know some of the side lengths of a right triangle? 4. What makes a triangle a “special” right triangle? How can special right triangles help me find side lengths? (Sorry to those who posted on my last question but the site mysteriously deleted it :[ )

OpenStudy (whpalmer4):

1. Pythagorean theorem says that the square of the length of the hypotenuse in a right triangle is the sum of the squares of the other two sides

OpenStudy (anonymous):

I mostly need help on number 3

OpenStudy (whpalmer4):

How many is "some"? :-)

OpenStudy (anonymous):

i assume 2

OpenStudy (anonymous):

i have this but i need the missing angle

OpenStudy (whpalmer4):

If you know 2 sides, then trigonometry should give you angles and remaining side.

OpenStudy (anonymous):

1. It matters if that triangle is a equilateral triangle or just a normal triangle... 2. Yes you can, you would have to use sin,cos and tang. 3. Yes you can, you would have to again use sin,cos and tang. 4. If all the angles are the same measurement then its special...It basically says that all the sides are the same length since they all have the same angle measurement.

OpenStudy (whpalmer4):

@zerosniper123 it said that it is a right triangle

OpenStudy (anonymous):

but what equation o use

OpenStudy (anonymous):

Ohh okay then you would use the Pythagorean theorem. :) And the formula is..... A^2+B^2=C^2

OpenStudy (anonymous):

i need the angle not length

OpenStudy (whpalmer4):

a special right triangle can be angle based (for example, 45 45 90 or 30 60 90) or it can be side based (3 4 5) there's a nice article on special right triangles on wikipedia

OpenStudy (anonymous):

thanks

OpenStudy (whpalmer4):

If you find BC with Pythagorean theorem, you know all the sides and can find all the angles

OpenStudy (whpalmer4):

Also, if you already know 1 angle + right angle, you know that the sum of the interior angles of a triangle = 180 degrees

OpenStudy (whpalmer4):

in general, a regular polygon with n sides the interior angles sum to (n-2)180 degrees. square is a regular polygon with 4 sides, so the interior angles sum to (4-2)180 = 360. we also know that a square has 4 right angles = 4 * 90 = 360.

OpenStudy (anonymous):

If you have atleast 2 sides, you should be able to figure the rest of it out using SOH CAH TOA